OpenStudy (anonymous):

Write the following expression using a single exponent. (8^4)^3 × 8^6

OpenStudy (anonymous):

(a^b)^c=a^(b*c) hence the given expression can be written as (8^(4*3))*(8^6)=8^(12+6)=8^18 @William14